UT Apr 07 2018 - Split Night - 1st half SOUL Testing / 2nd half LBTO Technical-Engineering

OSA: D.G. Huerta

ISp: J. Morris

Mtn Manager: R.Hansen

ISA: B. Rothberg (Tucson)
AO: D.Miller

Featured Players: J. Christou and the ARCETRI AO team

UT 03:13 - have been closed since sunset due to high humidity - last ~10 minutes its dropped below 90% - currently at 72% - will open if below 90% for 30min

UT 03:15 - having issue with ARGO Cal Unit being stuck - seems unable to be stowed correctly - G. Rahmer was called (message left)

UT 03:34 - planning to open enclosure now - cal unit issue fixed (thanks to G. Rahmer)

UT 03:45 - encloosure still closed - humidity ~72% - going to wait another 30min due to clouds

UT 04:01 - humidity just hit ~90%

UT 04:12- humidity at ~82% - still closed

UT 04:37 - humidity at 99.5%

UT 04:50 - humidity dropped again 83.8%

UT 04:53 - running closed dome cals with ARGOS cal lamp - however, can't find the light source - is M3 pointed at LBTI?

UT 04:54 - humidity is 61.6%

UT 05:27 - humidity 73.5% - all-sky fuzzy (condensation on camera or foggy outside?) - still closed

UT 05:47 - humidity ~74% - still closed

UT 06:03 - humidity ~67% - stlll closed - is all-sky foggy?

UT 06:17 - humidity 68.7% - still closed

UT 06:28 - Julian called up - some clouds still up there - not opening

UT 06:50 - humidity 77% - still closed

UT 07:28 - humidity 80.6% - ARGOS swing-arm in park position, but switch is not reading as in park - Doug working on it

UT 07:39 - ARCETRI team left for night

ENG/TECHNIAL PORTION OF NIGHT

UT 07:46 - ready to open - but need to call G. Rahmer since ARGOS Cal Unit won't park

UT 07:51 - bringing up LUCI GUIs - will do init all on LUCI-2

UT 08:06 - still having ARGOS calunit issues

L1.20190407.0007-0008 - 2.51 sec dark

L2.20190407.0001-0002 - 2.51 sec dark

UT 08:26 - G. Rahmer and D. Miller still working on ARGOS calunit issue

-will need to have Dan Cox remvoe PLC flag to override software thinking its not in park position

UT 08:38 - and we hit 90% humidity again

UT 08:49 - override in place, stowpins can be moved - humidity at 93%

-tipping telescope to elevation 70 clicked it and calunit says in park position

UT 09:23 - humidity 89.9%

UT 09:48 - humidity 87.8% - stil closed

UT 10:01 - humidity at 90.3%

UT 10:35 - humidity 87.2% - still closed

UT 10:45 - humidity 85% - still closed

UT 11:03 - humidity 86.7% - still closed

UT 11:17 - humidity 86%

UT 11:49- hit 95% humidity - into twilight

UT 12:00 - into nautical twilight - closing out all the windows - guess the night is done - no data taken

-- BarryRothberg - 07 Apr 2019
Topic revision: r1 - 07 Apr 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback